OpenStudy (anonymous):

Choose the correct domain of 5x^2+1/2x-1. the answer is All real numbers except 1/2 I NEED TO KNOW WHY =(

OpenStudy (amistre64):

when x=1/2; what do you get?

OpenStudy (anonymous):

0.5?

OpenStudy (amistre64):

no, do the math

OpenStudy (anonymous):

how did u come upon that answer though? what necessary steps would i have to take to stumble on that logic?

OpenStudy (amistre64):

what do you get when x=1/2 ?

OpenStudy (anonymous):

I'm not understanding your question? Do you mean when I plug it into the equation?

OpenStudy (amistre64):

yes, since there isnt really much else to do with it is there?

OpenStudy (anonymous):

u would get all decimals

OpenStudy (amistre64):

show me, do the math

OpenStudy (anonymous):

too long

OpenStudy (amistre64):

too long? then you really dont want to know the answer do you :/

OpenStudy (amistre64):

you get:\[\frac{2.5}{0}\]but that might take hours and days to figure out if we do it by hand right ....

OpenStudy (amistre64):

so, what does 2.5/0 equal?

OpenStudy (anonymous):

unknown

OpenStudy (anonymous):

or undefined

OpenStudy (amistre64):

correct, so if the answer cannot be determined; then we need to exclude it from the situation

OpenStudy (amistre64):

the domain is all real values that make sense; x=1/2 doesnt make sense to us

OpenStudy (anonymous):

oh ok i kinda get it thanx

OpenStudy (amistre64):

when the value of a variable makes a denominator go zero, we exclude it from the domain